Solve this question:
( 6 – x ) / ( x^2 – 4 ) = 2 + ( x /  (x + 2 )  ) )

Grade:11

1 Answers

Vijay Mukati
7 years ago
In LHS write x^2 – 4 = (x-2)(x+2) and take the LCM in RHS side only
Then you can easily cancel (x+2) factor on both sides of equaiton,
then simply cross multiply and simplify you should get the following quadratic equation:
3x^2 – x – 2 = 0, Now factorize it and get the two values of x from here.’

Thanks, Vj